# LMAct: A Benchmark for In-Context Imitation Learning with Long Multimodal Demonstrations

### v2

@@ Abstract
In this paper, we present a benchmark to pressure-test today’s frontier models’ multimodal decision-making capabilities in the very long-context regime (up to one million tokens) and investigate whether these models can learn from large numbers of expert demonstrations in their context. We evaluate the performance of Claude 3.5 Sonnet, Gemini 1.5 Flash, Gemini 1.5 Pro, Gemini 2.0 Flash Experimental, GPT-4o, o1-mini, o1-preview, and o1 as policies across a battery of simple interactive decision-making tasks: playing tic-tac-toe, chess, and Atari, navigating grid worlds, solving crosswords, and controlling a simulated cheetah. We study increasing amounts of expert demonstrations in the context — from no demonstrations to 512 full episodes. Across our tasks, models rarely manage to fully reach expert performance, and often presenting more demonstrations has little effect. Some models steadily improve with more demonstrations on a few tasks. We investigate the effect of encoding observations as text or images and the impact of chain-of-thought prompting. To help quantify the impact of other approaches and future innovations, we open source our benchmark that covers the zero-, few-, and many-shot regimes in a unified evaluation.
@@ 1 Introduction
While LMs have been shown to be able to perform non-trivial reasoning and decision-making in some domains (Huang et al., 2022; Yao et al., 2023; Romera-Paredes et al., 2024), there are also many negative results where LMs fail to perform decision-making tasks that are very simple for humans, even when LMs arguably possess great factual knowledge of the task. For example, LMs struggle to play legal moves, let alone beat amateur humans, in chess (Carlini, 2023). At the same time, state-of-the-art LMs have detailed expert knowledge of chess when queried in natural language. But this declarative knowledge fails to translate into effective decision-making, i.e., “know-how” (Ryle, 1949). This “knowing-doing gap” is also observed in the recently released BALROG benchmark (Paglieri et al., 2024), which evaluates zero-shot capabilities (i.e., without expert demonstrations) of state-of-the-art LMs on interactive decision-making tasks in the long time horizon setting: 5 game environments from Baby AI (Chevalier-Boisvert et al., 2019) to Nethack (Küttler et al., 2020). Overall, Paglieri et al. (2024) find that “models struggle significantly with more challenging tasks” and highlight few- and many-shot evaluation as an important potential solution. Our benchmark (developed concurrently) addresses exactly this gap by covering the full range from zero-shot to many-shot evaluation.
@@ 3.2 In-Context Imitation Learning
DM Control – Cheetah Run
@@ 4 Discussion & Related Work
Many large-scale benchmarks have been developed to test the general capabilities of frontier LMs, including the Chatbot Arena (Chiang et al., 2024) or LiveBench (White et al., 2024). Other benchmarks, such as FrontierMath (Glazer et al., 2024), TaskBench (Shen et al., 2023), GameBench (Costarelli et al., 2024), and Atari-GPT (Waytowich et al., 2024), specifically investigate the reasoning or interactive decision-making capabilities of (L)LMs. Most closely related to our work, and published in parallel, the BALROG (Paglieri et al., 2024) benchmark evaluates frontier LMs’ zero-shot reasoning on multimodal decision-making tasks in five game environments from Baby AI to NetHack. Our work complements these benchmarks by focusing on in-context imitation learning with long multimodal context, filling a gap in the current literature. Our benchmark is currently in its “easiest” form (e.g., grid world without obstacles) and can easily be made more challenging.
@@
We perform an evaluation via closed-source APIs and thus have little control over how the data is processed and fed to the underlying models. Since models behind the APIs can be updated at any time, it is possible that our results may not be quantitatively reproducible soon after publishing this manuscript. Despite our best efforts in evaluating different prompt formats, we cannot rule out that even small changes to the prompt could lead to better results. Accordingly, our current results are a lower bound on the models’ performance. Finally, we cannot guarantee that our observation formats do not cause tokenization issues across all APIs (e.g., a loss of structure for a 2D grid in ASCII).
@@ Acknowledgments
We thank Anna Mitenkova, Dipanjan Das, Jordi Grau-Moya, Joel Veness, Kate Olszewska, Li Kevin Wenliang, Laurent Orseau, Marcus Hutter, Matthew Aitchison, Matthew Ginsberg, Minmin Chen, Orhan Firat, Satinder Baveja, Shaobo Hou, and Zhengdong Wang, for their helpful feedback and insightful discussions.
@@ Appendix A Related Work
Instead of pretraining separate models for general decision-making, many researchers have also attempted to directly use the knowledge and reasoning capabilities of LLMs and VLMs for building agents that interact with an environment (see Xi et al. (2023) for a 2023 survey of LLM-based agents) — either by fine-tuning (Li et al., 2022) or purely via in-context demonstrations (Palo & Johns, 2024). In both cases, two open question are: (i) how to best represent environment observations as tokens, and (ii) how to best elicit the decision-making capabilities of pretrained LMs. Mirchandani et al. (2023) find that pretrained LLMs are “general pattern machines” that can learn to complete complex token sequences via in-context learning, including agent-environment interactions. They even find that in many cases, randomly swapping the alphabet does not have significant impact, suggesting that LLMs may be able to deal with many different ways of translating observations into tokens. Perhaps more important is the question whether observations should be state-action sequences (as in imitation learning and our work) or whether they should include rewards (for reward conditioning or in-context RL as in Mirchandani et al. (2023) and Raparthy et al. (2024)). While this is still unclear for pretrained LMs, Ruoss et al. (2024) find that, when training a large transformer to play chess, performance is roughly the same for imitation learning compared to learning to predict state or action values (as long as the amount of data for all three variants is equal). Schultz et al. (2024) extend these results on chess to LLMs by distilling the search proceedure into the model by linearizing the search trees. Ma et al. (2024) find that, when prompted appropriately, a pretrained VLM (Gemini 1.5 Pro) can produce good value estimates for real-world robotic tasks. The second open problem, how to best prompt LMs for decision-making, is a very active research area, with a lot of focus on designing or learning zero-shot prompts, such as the famous “Let’s think about this step by step.” (Kojima et al., 2022) and “Take a deep breath and work on this problem step-by-step.” (Yang et al., 2024), which has led to many chain-of-thought prompting schemes (Wei et al., 2022). Besides better zero-shot prompts, advanced sampling and prompt optimization schemes have been explored, such as iterative prompt refinement where an agent starts with some demonstrations in the context, then interacts with the environment, and potentially replaces a lower performing episode with a higher performing one (Mirchandani et al., 2023; Brooks et al., 2023).
@@
With the recent availability of long context models, a third possibility compared to improving zero-shot or few-shot prompts has emerged: many-shot in-context learning, i.e., prompting with many demonstrations, on the order of having a full small dataset in the context. Both Agarwal et al. (2024) and Jiang et al. (2024) show that many-shot prompts (hundreds or thousands of examples in the prompt) improve pretrained LM performance over few-shot prompts in non-interactive tasks. Our paper explores the same direction, with potentially hundreds of full episodes in the context, for interactive decision-making tasks. At the time of writing, querying long context models with many tokens comes with high computational cost. An alternative could be to virtually extend the context via retrieval based methods, e.g., REGENT (Sridhar et al., 2024), which trains a retrieval based agent. The retrieval problem is currently not fully solved (as also pointed out in Paglieri et al. (2024)), and today’s largest state-of-the-art LMs do not offer retrieval via their APIs and/or have not been trained to perform retrieval. Accordingly, placing all the data in the context, as in our work, allows estimating LM performance independent of whether retrieval works well. Our results thus serve as an upper baseline to calibrate retrieval based methods against.
@@
Comparing capabilities of LMs that span many tasks and domains requires large-scale benchmarks, such as the Chatbot Arena (a.k.a. LMSys, Chiang et al. (2024)) or LiveBench (White et al., 2024). Other benchmarks, such as FrontierMath (Glazer et al., 2024), Taskbench (Shen et al., 2023), and Gamebench (Costarelli et al., 2024), specifically investigate the reasoning or interactive decision-making capabilities of LLMs. Closely related to these and to our work, and published in parallel to our work, the BALROG (Paglieri et al., 2024) benchmark evaluates state-of-the-art LMs (the same as in our work with the exception of o1-mini and o1-preview and the addition of the Llama 3 models, which we do not evaluate) on multimodal reasoning and decision-making tasks by using a set of 5 increasingly harder game environments from Baby AI to Nethack. Like our work, the authors find that state-of-the-art LMs struggle significantly in challenging game environments. The authors also make the observation that models arguably possess a lot of knowledge about their tasks when queried appropriately, but that state-of-the-art LMs have a “knowing-doing” gap. Unlike our work, the BALROG paper performs zero-shot evaluation and lists few-shot and many-shot evaluations as an important open research question (the released codebase supports few-shot evaluations, but these evaluations have not been performed at the time of writing). Similar to BALROG, Waytowich et al. (2024) evaluate the zero-shot game-playing capabilities of frontier LLMs (GPT-4V Turbo, GPT-4o, Gemini 1.5 Flash, and Claude $3$ Haiku) on $8$ different Atari games (but not the Phoenix game that we consider). Compared to these previous benchmarks, ours is the only one that puts the emphasis on imitation learning with long context in multimodal interactive environments — a regime that pushes against the limits of modern LMs’ in-context learning and reasoning capabilities, both from an engineering and a capabilities perspective. Furthermore, our benchmark covers the zero-shot, few-shot, and many-shot setting in a unified and thus easily comparable evaluation.
@@ C.3 Illegal Actions
As described in Section 2.4, we do not terminate the evaluation episode early if a model does not produce a valid action (except on the crossword task) but instead randomly sample one of the actions that are legal in the current state and continue the evaluation with the observation produced by that action. To differentiate illegal actions from acting randomly (over legal actions), we compute how often models actually propose an illegal action and visualize the percentage of illegal actions per episode over the number of expert demonstrations in Figs. A15, A16, A17, A18, A19 and A20 (analogous to Figs. 3, 4, 5, 6, 7 and 8). Note that, since we perform an ablation over showing the legal moves in the prompt (see Section C.4), in theory, all models could have the necessary information to sample a legal action at all times (whether we actually show this or not in our main experiments depends on whether it improves model performance in the ablations). Nevertheless, we observe that models do produce illegal actions actions across most environments. For example, on the grid world navigation task with coordinate observations o1-preview produces almost $100$% illegal actions with $2$ or more expert demonstrations in the context. Interestingly, the ability to produce legal actions also depends on the observation type: For tic-tac-toe, all models mostly produce legal actions with ASCII observations, but the Gemini 1.5 models generate up to $\sim 60$% illegal actions for RGB observations from the same environment.

### v3

@@ Abstract
In this paper, we present a benchmark to pressure-test today’s frontier models’ multimodal decision-making capabilities in the very long-context regime (up to one million tokens) and investigate whether these models can learn from large numbers of expert demonstrations in their context. We evaluate the performance of Claude 3.5 Sonnet, Gemini 1.5 Flash, Gemini 1.5 Pro, Gemini 2.0 Flash Experimental, GPT-4o, o1-mini, o1-preview, and o1 as policies across a battery of simple interactive decision-making tasks: playing tic-tac-toe, chess, and Atari, navigating grid worlds, solving crosswords, and controlling a simulated cheetah. We study increasing amounts of expert demonstrations in the context — from no demonstrations to 512 full episodes. Across our tasks, models rarely manage to fully reach expert performance, and often, presenting more demonstrations has little effect. Some models steadily improve with more demonstrations on a few tasks. We investigate the effect of encoding observations as text or images and the impact of chain-of-thought prompting. To help quantify the impact of other approaches and future innovations, we open source our benchmark that covers the zero-, few-, and many-shot regimes in a unified evaluation.
@@ 1 Introduction
While LMs have been shown to be able to perform non-trivial reasoning and decision-making in some domains (Huang et al., 2022; Yao et al., 2023; Romera-Paredes et al., 2024), there are also many negative results where LMs fail to perform decision-making tasks that are very simple for humans, even when LMs arguably possess great factual knowledge of the task. For example, LMs struggle to play legal moves, let alone beat amateur humans, in chess (Carlini, 2023). At the same time, state-of-the-art LMs have detailed expert knowledge of chess when queried in natural language. But this declarative knowledge fails to translate into effective decision-making, i.e., “know-how” (Ryle, 1949). This “knowing-doing gap” is also observed in the recently released BALROG benchmark (Paglieri et al., 2025), which evaluates zero-shot capabilities (i.e., without expert demonstrations) of state-of-the-art LMs on interactive decision-making tasks in the long time horizon setting: 5 game environments from Baby AI (Chevalier-Boisvert et al., 2019) to Nethack (Küttler et al., 2020). Overall, Paglieri et al. (2025) find that “models struggle significantly with more challenging tasks” and highlight few- and many-shot evaluation as an important potential solution. Our benchmark (developed concurrently) addresses exactly this gap by covering the full range from zero-shot to many-shot evaluation.
@@ 3.2 In-Context Imitation Learning
Chess
Tic-Tac-Toe
(a) ASCII observations
(b) RGB observations
@@ 4 Discussion & Related Work
Many large-scale benchmarks have been developed to test the general capabilities of frontier LMs, including the Chatbot Arena (Chiang et al., 2024) or LiveBench (White et al., 2025). Other benchmarks, such as FrontierMath (Glazer et al., 2024), TaskBench (Shen et al., 2024), GameBench (Costarelli et al., 2024), and Atari-GPT (Waytowich et al., 2024), specifically investigate the reasoning or interactive decision-making capabilities of (L)LMs. Most closely related to our work, and published in parallel, the BALROG (Paglieri et al., 2025) benchmark evaluates frontier LMs’ zero-shot reasoning on multimodal decision-making tasks in five game environments from Baby AI to NetHack. Our work complements these benchmarks by focusing on in-context imitation learning with long multimodal context, filling a gap in the current literature. Our benchmark is currently in its “easiest” form (e.g., grid world without obstacles) and can easily be made more challenging.
@@
We perform an evaluation via closed-source APIs and thus have little control over how the data is processed and fed to the underlying models. Since models behind the APIs can be updated at any time, it is possible that our results may not be quantitatively reproducible soon after publishing this manuscript. Despite our best efforts in evaluating different prompt formats, we cannot rule out that even small changes to the prompt could lead to better results. Accordingly, our current results are a lower bound on the models’ performance. We can also not guarantee that our observation formats do not cause tokenization issues across all APIs (e.g., a loss of structure for a 2D grid in ASCII). Finally, while our environments are simpler than complex real-world scenarios like robotics, they require the same fundamental agentic capabilities (long-context multimodal understanding, in-context imitation learning) that are often tested in robotics research and likely necessary for real-world success. Accordingly, our LMAct benchmark serves as a controlled testbed to evaluate these core skills, diagnose current model limitations, and guide research toward more generally capable agents. Although direct transfer is beyond this paper’s scope, findings on our benchmark can inform future work addressing that challenge.
Future Work
For Cheetah Run (Fig. 8) and grid world navigation with ASCII demonstrations (Fig. 7) we observe that the models’ performance deteriorates with increasing numbers of expert demonstrations, a phenomenon we refer to as “in-context interference”. For both tasks our analysis of the percentage of illegal actions (Figs. A18 and A19) suggests the problem is not due to an increased number of illegal actions, but, instead, due to increasingly suboptimal actions. We have conducted a set of initial investigations into these failure modes (see Sections C.2, C.3, C.4 and A14), but a definitive answer to what causes in-context interference would require a thorough investigation, and, therefore, presents an interesting direction for future work. Another promising avenue for future work is to focus on models capable of general in-context reinforcement learning, which is a bit different than our in-context imitation setting (in principle, all our tasks could easily be extended by providing additional reward observations). It also seems plausible that pretraining or finetuning with data from interactive decision-making tasks, and, in particular, in-context imitation of an expert policy, would be quite effective. Finally, we also think that evaluating LMs’ performance on partially observable tasks presents an interesting direction for future research (we primarily investigate fully observable tasks). Partially observable tasks require consistent integration of information across several, potentially non-adjacent, time steps, which is a great test of a model’s ability to densely attend to the information in the context. However, under partial observability, there is a theoretical problem of self-delusion in imitation of an expert that has hidden information (e.g., the hidden belief state of the agent; see Ortega et al. (2021)). We want to keep our benchmark free from these complicating issues and think that the right time to move to such harder tasks is when frontier models easily solve simple, fully observable tasks at expert level (we are currently at beginner level). Nevertheless, we believe there is great value in benchmarks on interactive decision-making tasks under partial observability (perhaps better suited for in-context reinforcement learning, which avoids the self-delusion problem), and refer to, e.g., the BALROG benchmark (Paglieri et al., 2025).
@@ Acknowledgments
We thank Anna Mitenkova, Dipanjan Das, Jordi Grau-Moya, Joel Veness, Kate Olszewska, Li Kevin Wenliang, Laurent Orseau, Marcus Hutter, Matthew Aitchison, Matthew Ginsberg, Minmin Chen, Orhan Firat, Satinder Baveja, Shaobo Hou, Zhengdong Wang, and the anonymous reviewers for their helpful feedback and insightful discussions.
@@ Appendix A Related Work
Instead of pretraining separate models for general decision-making, many researchers have also attempted to directly use the knowledge and reasoning capabilities of LLMs and VLMs for building agents that interact with an environment (see Xi et al. (2025) for a 2023 survey of LLM-based agents) — either by fine-tuning (Li et al., 2022) or purely via in-context demonstrations (Palo & Johns, 2024). In both cases, two open question are: (i) how to best represent environment observations as tokens, and (ii) how to best elicit the decision-making capabilities of pretrained LMs. Mirchandani et al. (2023) find that pretrained LLMs are “general pattern machines” that can learn to complete complex token sequences via in-context learning, including agent-environment interactions. They even find that in many cases, randomly swapping the alphabet does not have significant impact, suggesting that LLMs may be able to deal with many different ways of translating observations into tokens. Perhaps more important is the question whether observations should be state-action sequences (as in imitation learning and our work) or whether they should include rewards (for reward conditioning or in-context RL as in Mirchandani et al. (2023) and Raparthy et al. (2024)). While this is still unclear for pretrained LMs, Ruoss et al. (2024) find that, when training a large transformer to play chess, performance is roughly the same for imitation learning compared to learning to predict state or action values (as long as the amount of data for all three variants is equal). Schultz et al. (2025) extend these results on chess to LLMs by distilling the search proceedure into the model by linearizing the search trees. Ma et al. (2025) find that, when prompted appropriately, a pretrained VLM (Gemini 1.5 Pro) can produce good value estimates for real-world robotic tasks. The second open problem, how to best prompt LMs for decision-making, is a very active research area (Wenliang et al., 2025; Genewein et al., 2025), with a lot of focus on designing or learning zero-shot prompts, such as the famous “Let’s think about this step by step.” (Kojima et al., 2022) and “Take a deep breath and work on this problem step-by-step.” (Yang et al., 2024), which has led to many chain-of-thought prompting schemes (Wei et al., 2022). Besides better zero-shot prompts, advanced sampling and prompt optimization schemes have been explored, such as iterative prompt refinement where an agent starts with some demonstrations in the context, then interacts with the environment, and potentially replaces a lower performing episode with a higher performing one (Mirchandani et al., 2023; Brooks et al., 2023).
@@
With the recent availability of long context models, a third possibility compared to improving zero-shot or few-shot prompts has emerged: many-shot in-context learning, i.e., prompting with many demonstrations, on the order of having a full small dataset in the context. Both Agarwal et al. (2024) and Jiang et al. (2024) show that many-shot prompts (hundreds or thousands of examples in the prompt) improve pretrained LM performance over few-shot prompts in non-interactive tasks. Our paper explores the same direction, with potentially hundreds of full episodes in the context, for interactive decision-making tasks. At the time of writing, querying long context models with many tokens comes with high computational cost, but it would be interesting to investigate how specialized models that were specifically developed for long-context tasks (Bulatov et al., 2022; Cherepanov et al., 2024) would fare on our benchmark. An alternative could be to virtually extend the context via retrieval based methods, e.g., REGENT (Sridhar et al., 2024), which trains a retrieval based agent. The retrieval problem is currently not fully solved (as also pointed out in Paglieri et al. (2025)), and today’s largest state-of-the-art LMs do not offer retrieval via their APIs and/or have not been trained to perform retrieval. Accordingly, placing all the data in the context, as in our work, allows estimating LM performance independent of whether retrieval works well. Our results thus serve as an upper baseline to calibrate retrieval based methods against.
Another line of work has investigated in-context learning for control (Duan et al., 2017; Xu et al., 2022; Fu et al., 2024; Fang et al., 2024; Yin et al., 2024; Wang et al., 2024). While these methods focus on developing novel agents with strong in-context learning capabilities, our paper’s primary goal is to benchmark the current state of existing, general-purpose frontier LMs on such tasks. Understanding how to best leverage insights from specialized agent research to enhance these large, pre-trained models remains an open question, which underscores the importance of establishing clear benchmarks like LMAct.
@@
Comparing capabilities of LMs that span many tasks and domains requires large-scale benchmarks, such as the Chatbot Arena (a.k.a. LMSys, Chiang et al. (2024)) or LiveBench (White et al., 2025). Other benchmarks, such as FrontierMath (Glazer et al., 2024), Taskbench (Shen et al., 2024), Gamebench (Costarelli et al., 2024), and BABILong (Kuratov et al., 2024), specifically investigate the reasoning or interactive decision-making capabilities of LLMs. Closely related to these and to our work, and published in parallel to our work, the BALROG (Paglieri et al., 2025) benchmark evaluates state-of-the-art LMs (the same as in our work with the exception of o1-mini and o1-preview and the addition of the Llama 3 models, which we do not evaluate) on multimodal reasoning and decision-making tasks by using a set of 5 increasingly harder game environments from Baby AI to Nethack. Like our work, the authors find that state-of-the-art LMs struggle significantly in challenging game environments. The authors also make the observation that models arguably possess a lot of knowledge about their tasks when queried appropriately, but that state-of-the-art LMs have a “knowing-doing” gap. Unlike our work, the BALROG paper performs zero-shot evaluation and lists few-shot and many-shot evaluations as an important open research question (the released codebase supports few-shot evaluations, but these evaluations have not been performed at the time of writing). Similar to BALROG, Waytowich et al. (2024) evaluate the zero-shot game-playing capabilities of frontier LLMs (GPT-4V Turbo, GPT-4o, Gemini 1.5 Flash, and Claude $3$ Haiku) on $8$ different Atari games (but not the Phoenix game that we consider). Compared to these previous benchmarks, ours is the only one that puts the emphasis on imitation learning with long context in multimodal interactive environments — a regime that pushes against the limits of modern LMs’ in-context learning and reasoning capabilities, both from an engineering and a capabilities perspective. Furthermore, our benchmark covers the zero-shot, few-shot, and many-shot setting in a unified and thus easily comparable evaluation.
@@ C.3 Illegal Actions
As described in Section 2.4, we do not terminate the evaluation episode early if a model does not produce a valid action (except on the crossword task) but instead randomly sample one of the actions that are legal in the current state and continue the evaluation with the observation produced by that action. To differentiate illegal actions from acting randomly (over legal actions), we compute how often models actually propose an illegal action and visualize the percentage of illegal actions per episode over the number of expert demonstrations in Figs. A15, A16, A17, A18, A19 and A20 (analogous to Figs. 3, 4, 5, 8, 7 and 6). Note that, since we perform an ablation over showing the legal moves in the prompt (see Section C.4), in theory, all models could have the necessary information to sample a legal action at all times (whether we actually show this or not in our main experiments depends on whether it improves model performance in the ablations). Nevertheless, we observe that models do produce illegal actions actions across most environments. For example, on the grid world navigation task with coordinate observations o1-preview produces almost $100$% illegal actions with $2$ or more expert demonstrations in the context. Interestingly, the ability to produce legal actions also depends on the observation type: For tic-tac-toe, all models mostly produce legal actions with ASCII observations, but the Gemini 1.5 models generate up to $\sim 60$% illegal actions for RGB observations from the same environment.